OpenStudy (anonymous):

\[\int\limits \frac{dx}{x \sqrt{x ^{2}-16}}\]

OpenStudy (foolaroundmath):

\(x = 4\sec{\theta} \implies \text{d}x = 4\sec{\theta}\tan{\theta}\text{d}\theta\)

OpenStudy (anonymous):

You are looking for a substitution that will help you to get rid of that radical, let me show you why @FoolAroundMath substitution works. If \(x=4\sec(\theta) \) then the radical becomes \[ 4\sqrt{\sec^2(\theta) -1}=4 \tan(\theta)\]